OpenStudy (anonymous):

The sum of two numbers is 170 and their difference is 26. Find the two numbers. Show how to solve this using any method.

OpenStudy (anonymous):

How would you solve this?

OpenStudy (anonymous):

Would the 2 numbers be 98 and -72 ?

OpenStudy (anonymous):

x+y=170, y-x=26 y-x=26 y=26+x x+26+x=170 2x+26=170 2x=170-26 2x=144 x=72 x+y=170 72+y=170 y=170-72 y=98.

OpenStudy (anonymous):

good patty...

OpenStudy (anonymous):

hehehehe.. over2x.. iclick na gen ang naa sa Good answer para naa q medal.

OpenStudy (gw2011):

The numbers would be 98 and 72. You can solve this by using simultaneous equations: x + y = 170 x - y = 26 Now add the two equations to get: 2x = 196 x = 98 Now substitute x=98 into one of the equations to get the other number--for example, x + y = 170 98 + y = 170 y = 72